Circular No. 5718

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993G IN NGC 3690 R. R. Treffers, B. Leibundgut, and A. V. Filippenko, University of

California at Berkeley; and M. W. Richmond, Princeton University, report their discovery of an apparent supernova in NGC 3690 (α = 11^{h}25^{m}44^{s}.2, δ = +58^{o}50′23′′, equinox 1950.0). Observations were conducted as part of the Leuschner Observatory Supernova Search (LOSS), which uses an automated 0.76-m telescope equipped with the Lawrence Berkeley Laboratory CCD camera. The supernova is located ∼ 1′′.5 west and 15′′ south of the galaxy’s eastern, relatively diffuse, condensation. The discovery image was taken on Mar. 5 UT, with confirmation on Mar. 6. The supernova is not present in previous LOSS images of the galaxy, the most recent one having been obtained on Feb. 15. Data obtained on Mar. 6.26 with the 0.5-m Berkeley Automatic Imaging Telescope (BAIT) yield R = 16.6 ± 0.2 for the supernova.

V1055 ORIONIS R. M. Wagner, D. L. DePoy, R. D. Blum, and R. Bertram, Ohio State

University; and D. T. Thompson and M. W. Buie, Lowell Observatory, report: “V - and K-band imagery of V1055 Ori (IAUC 5717) were obtained on Mar. 5.1 UT at the Hall 1.1-m and Perkins 1.8-m telescopes at the Lowell Observatory using the Lowell RCA direct CCD camera and the newlycommissioned Ohio State Infrared Imaging System, respectively. Digital aperture photometry of these data gives V = 20.0 ± 0.7 and K = 17.2 ± 0.3, with the large error in the V magnitude attributable to the small telescope and a bright background from the nearly full moon. Our results are consistent with previous photometry of V1055 Ori (Machin et al. 1990, MNRAS 247, 205; Gottwald et al. 1991, A.Ap. Suppl. 89, 367), which give a mean V = 18.9 ± 0.1, but with some intrinsic variability of ∼ 0.2-0.3 mag. Our photometry indicates that V1055 Ori appears to be near its quiescent brightness and not in an unusual high state, but the observations cover only one epoch. Additional photometric observations are encouraged.”

CY URSAE MAJORIS Visual magnitude estimates by P. Schmeer, Bischmisheim, Germany,

of a superoutburst of this SU UMa-type dwarf nova: Feb. 23.00 UT, [13.2; 24.07, 12.2; 24.84, 12.3; 25.88, 12.5; 28.20, 12.7; Mar. 1.11, 12.7; 3.15, 12.9.

1993 March 6 Daniel W. E. Green